1. Find the rotational kinetic energy of an engine flywheel that revolve at 30 rad/s if the moment of inertia is 4,000 kg.m2
1
Expert's answer
2021-03-14T19:14:25-0400

"E=\frac{I\omega^2}{2}=\frac{4000\cdot30^2}{2}=1800000\ (J)=1.8\ (MJ)" . Answer